### v1

# A Recipe For Obtaining Algebraic Addition Theorems Of The Weierstrass Elliptic Function

Abstract
In this paper, we present a general method for obtaining addition theorems of the Weierstrass elliptic function $\wp(z)$ in terms of given parameters. We obtain the classical addition theorem for the Weierstrass elliptic function as a special case. Furthermore, we give novel two-term addition, three-term addition, duplication and triplication formulas. New identities for elliptic invariants are also proven.
Keywords: Elliptic functions, Weierstrass elliptic function, Addition theorems, Elliptic invariants
AMS Subject Classifications: 33E05
1 Introduction
Let $\Omega=(2\omega_{1},2\omega_{2})\in\mathbb{C}^{2}$ be such that $\text{Im}(\omega_{2}/\omega_{1})>0$ and $\Lambda_{\Omega}=2\omega_{1}\mathbb{Z}+2\omega_{2}\mathbb{Z}$. The Weierstrass elliptic function Weierstrass-$\wp$ is defined as
$\wp(z)=\wp(z;\Lambda_{\Omega})=\frac{1}{z^{2}}+\sum_{\omega\in\Lambda_{\Omega}^{*}}\frac{1}{(z-\omega)^{2}}-\frac{1}{\omega^{2}},$
where $\Lambda_{\Omega}^{*}=\Lambda_{\Omega}-\{0\}$ and the series converges absolutely and uniformly except at the poles $z\in\Lambda_{\Omega}$. It is an even elliptic function of order $2$ with double poles at $z\in\Lambda_{\Omega}$. Weierstrassian elliptic invariants of the lattice $\Lambda_{\Omega}$ are defined as
$g_{2}=60\sum_{\omega\in\Lambda_{\Omega}^{*}}\omega^{-4}\qquad\text{and}\qquad g_{3}=140\sum_{\omega\in\Lambda_{\Omega}^{*}}\omega^{-6}.$
The function Weierstrass-$\wp$ satisfies the well-known algebraic differential equation
$\left(\wp^{\prime}(z)\right)^{2}=4\wp(z)^{3}-g_{2}\wp(z)-g_{3}=4(\wp(z)-e_{1})(\wp(z)-e_{2})(\wp(z)-e_{3}).$ (1)
Here $e_{1}=\wp(\omega_{1})$, $e_{2}=\wp(\omega_{2})$ and $e_{3}=\wp(\omega_{1}+\omega_{2})$ are the half-period values. We have the famous formulae
$e_{1}+e_{2}+e_{3}=0$
$e_{1}e_{2}+e_{2}e_{3}+e_{3}e_{1}=-g_{2}/4$
$e_{1}e_{2}e_{3}=g_{3}/4.$
Let $\sigma(z;\Lambda_{\Omega})$ denote the Weierstrass $\sigma$-function, defined by
$\sigma(z)=\sigma(z;\Lambda_{\Omega})=z\prod_{\omega\in\Lambda_{\Omega}^{*}}\left(1-\frac{z}{\omega}\right)\exp\left(\frac{z}{\omega}+\frac{z^{2}}{2\omega^{2}}\right).$
We have the basic relation $\left(\log\sigma(z)\right)^{\prime\prime}=-\wp(z)$. Algebraic addition theorems are fundamental objects in the theory of elliptic functions. Frobenius-Stickelberger determinant identity or pseudo-addition theorem [4, 9] states that
$\begin{vmatrix}1&\wp(z_{0})&\wp^{\prime}(z_{0})&\cdots&\wp^{(n-1)}(z_{0})\\ 1&\wp(z_{1})&\wp^{\prime}(z_{1})&\cdots&\wp^{(n-1)}(z_{1})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 1&\wp(z_{n})&\wp^{\prime}(z_{n})&\cdots&\wp^{(n-1)}(z_{n})\end{vmatrix}=(-1)^{\frac{n(n-1)}{2}}1!2!\cdots n!\frac{\sigma(z_{0}+z_{1}+\cdots+z_{n})}{\sigma(z_{0})^{n+1}\cdots\sigma(z_{n})^{n+1}}\prod_{0\leq i<j\leq n}\sigma(z_{i}-z_{j}),$ (2)
and for $n=1$ reduces to the well-known formula
$\wp(z)-\wp(w)=\frac{\sigma(w+z)\sigma(w-z)}{\sigma(w)^{2}\sigma(z)^{2}}.$
Frobenius-Stickelberger identities may be used to obtain algebraic addition theorems for the function Weierstrass-$\wp$ (for example, see [1, 8]). For more related determinant identities we refer the reader to [4, 5, 6, 7, 9]. The classical addition theorem for Weierstrass-$\wp$ is given as follows.
Theorem 1.1.
Let $z,w,z\pm w\notin\Lambda_{\Omega}$, then we have
$\begin{vmatrix}1&\wp(z)&\wp^{\prime}(z)\\ 1&\wp(w)&\wp^{\prime}(w)\\ 1&\wp(z+w)&-\wp^{\prime}(z+w)\end{vmatrix}=0$
Furthermore,
$\wp(z+w)=\frac{1}{4}\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}-\wp(z)-\wp(w).$
Upon taking the limit $w\to z$, we also obtain the duplication formula
$\wp(2z)=\frac{1}{4}\left(\frac{\wp^{\prime\prime}(z)}{\wp^{\prime}(z)}\right)^{2}-2\wp(z).$
For recent advances in addition theorems for elliptic functions, see [2, 3]. In this note, we discuss a recipe to obtain algebraic addition theorems for Weierstrass-$\wp$, expressing $\wp(z_{1}+\cdots+z_{\ell})$ in terms of $\wp^{(j)}(z_{i})$ for predetermined $j$.
2 Main Results
For convenience, let $\wp^{(-2)}(z)\equiv 1$ denote the constant function. Thus $\wp^{(i)}$ is an elliptic function of order $i+2$. Throughout this chapter, we assume that $\{n_{i}\}_{i=1}^{m}$, $\{k_{i}\}_{i=1}^{\ell}$ are integers such that $n_{i}\in\mathbb{N}\cup\{-2\}$ for every $i=1,\cdots,m$ and $k_{i}\in\mathbb{N}\cup\{-2\}$ for every $i=1,\cdots,\ell$. Assume $n_{i}\neq n_{j}$ for $i\neq j$ and $k_{i}\neq k_{j}$ for $i\neq j$. Let $\{\gamma_{i}\}_{i=1}^{m}$ and $\{z_{i}\}_{i=1}^{\ell}$ be given complex numbers such that $z_{i}\notin\Lambda_{\Omega}$ for every $i=1,\ldots,\ell$ and
$\begin{vmatrix}\wp^{(k_{1})}(z_{1})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \wp^{(k_{1})}(z_{2})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \wp^{(k_{1})}(z_{\ell})&\wp^{(k_{2})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\end{vmatrix}\neq 0.$ (3)
Note that this implies $z_{i}\not\equiv z_{j}\bmod\Lambda_{\Omega}$ for $i\neq j$. Let $n_{a}=\max n_{i}$, $k_{b}=\max k_{i}$ and $\ell=\max(n_{a},k_{b})+1$. Thus there exists complex numbers $\{\lambda_{i}\}_{i=1}^{\ell}$ such that the system
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{j})=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(z_{j})$ (4)
holds for every $j=1,\ldots,\ell$. Let us denote $z=z_{1}+\cdots+z_{\ell}$. Finally, assume that $z\notin\Lambda_{\Omega}$ and both of the equations $n_{a}=k_{b}$, $\gamma_{a}=\lambda_{b}$ do not hold at the same time.
We use Liouville’s theorems and elementary facts about elliptic functions frequently without mention. We first prove the determinant form of the addition theorem.
Theorem 2.1.
The following equation holds,
$\begin{vmatrix}\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\wp^{(k_{1})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\wp^{(k_{1})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\wp^{(k_{1})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)&\wp^{(k_{1})}(-z)&\cdots&\wp^{(k_{\ell})}(-z)\end{vmatrix}=0.$
Proof.
Let us consider the function
$\psi_{n,k,\gamma,\lambda}(s)=\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(s).$
By our assumptions, $\psi_{n,k,\gamma,\lambda}$ is an elliptic function of order $\max(n_{a},k_{b})+2=\ell+1$ with a pole of order $\ell+1$ at $s=0$. Since the equation (4) holds for any $j=1,\ldots,\ell$, the function $\psi_{n,k,\gamma,\lambda}$ has $\ell$ distinct zeros, namely $z_{1},\ldots,z_{\ell}$. Let $w$ denote the $\ell+1$th zero. Using Liouville’s theorem we obtain $w+\sum_{i=1}^{\ell}z_{i}\equiv 0\bmod\Lambda_{\Omega}$ and $w\equiv-z\bmod\Lambda_{\Omega}.$ Thus $-z$ is a root of $\psi_{n,k,\gamma,\lambda}$ and we get
$\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)=\sum_{i=1}^{\ell}\lambda_{i}\wp^{(k_{i})}(-z).$ (5)
Putting equations (4) and (5) together gives
$\begin{pmatrix}\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\wp^{(k_{1})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\wp^{(k_{1})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\wp^{(k_{1})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)&\wp^{(k_{1})}(-z)&\cdots&\wp^{(k_{\ell})}(-z)\end{pmatrix}\begin{pmatrix}-1\\ \lambda_{1}\\ \vdots\\ \lambda_{\ell}\end{pmatrix}=-\begin{pmatrix}\psi_{n,k,\gamma,\lambda}(z_{1})\\ \psi_{n,k,\gamma,\lambda}(z_{2})\\ \vdots\\ \psi_{n,k,\gamma,\lambda}(z_{\ell})\\ \psi_{n,k,\gamma,\lambda}(-z)\end{pmatrix}=0.$ (6)
Since the vector on the left-hand side of (6) is non-zero, the matrix on the left-hand side is singular. Therefore we have
$\begin{vmatrix}\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\wp^{(k_{1})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\wp^{(k_{1})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\wp^{(k_{1})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\\ \sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(-z)&\wp^{(k_{1})}(-z)&\cdots&\wp^{(k_{\ell})}(-z)\end{vmatrix}=0.$
This completes the proof. ∎
We remark that since $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and Weierstrass-$\wp$,$\wp^{(i)}$ satisfy an algebraic relationship for every $i$, Theorem 2.1 is indeed an algebraic addition theorem.
We need the following two lemmas.
Lemma 2.2.
For all $i\in\mathbb{N}$, $\wp^{(2n)}\in\mathbb{Q}[g_{2},g_{2}][\wp]$ and $\wp^{(2n+1)}\in\wp^{\prime}\mathbb{Q}[g_{2},g_{2}][\wp]$. Furthermore $\wp^{(2n)}$ is a polynomial in Weierstrass-$\wp$ of degree $n+1$ and $\wp^{(2n+1)}/\wp^{\prime}$ a polynomial in Weierstrass-$\wp$ of degree $n$.
Proof.
The proof follows immediately after induction and equation (1). ∎
Lemma 2.3.
The function defined as
$\varphi_{n,k,\gamma,\lambda}(s)=\left(\sum_{2\nmid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\nmid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}-\left(\sum_{2\mid n_{i}}\gamma_{i}\wp^{(n_{i})}(s)-\sum_{2\mid k_{i}}\lambda_{i}\wp^{(k_{i})}(s)\right)^{2}$ (7)
is a polynomial in $\wp(s)$ of degree at most $\ell+1$.
Proof.
By Lemma 2.2 and equation (1), the first sum in (7) is in $\left(\wp^{\prime}\mathbb{C}[\wp]\right)^{2}$ and also in $\mathbb{C}[\wp]$. Similarly the second sum in (7) is a polynomial in $\mathbb{C}[\wp]$. Therefore $\varphi_{n,k,\gamma,\lambda}\in\mathbb{C}[\wp]$. Again by Lemma 2.2 and (1), the first sum is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{2\nmid n_{i},k_{j}}(n_{i},k_{j})+2$ and second sum is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{2\mid n_{i},k_{j}}(n_{i},k_{j})+2$. Thus $\varphi_{n,k,\gamma,\lambda}$ is a polynomial in Weierstrass-$\wp$ of degree at most $\max_{i,j}(n_{i},k_{j})+2=\ell+1$. ∎
We now see that there exists constants $\mu_{n,k,\gamma,\lambda}(r)\in\mathbb{C}$, $r=0,1,\ldots,\ell+1$ such that
$\varphi_{n,k,\gamma,\lambda}(s)=\sum_{r=0}^{\ell+1}\mu_{n,k,\gamma,\lambda}(r)\wp(s)^{r}.$
Let $S_{r}(x_{1},\cdots x_{n})$, $k\leq n$ denote the elementary symmetric polynomials defined by
$S_{r}(x_{1},\ldots,x_{n})=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}x_{i_{1}}\ldots x_{i_{k}}.$
Furthermore, apply the convenience of notation
$S_{r}\wp(x_{1},\ldots,x_{n})=S_{r}\left(\wp(x_{1}),\ldots,\wp(x_{n})\right).$
Assuming that $\varphi_{n,k,\gamma,\lambda}\not\equiv 0$, we now prove the explicit form of the addition theorem.
Theorem 2.4.
Let $z\not\equiv\pm z_{i}\bmod\Lambda_{\Omega}$ and $z_{i}\not\equiv-z_{j}\bmod\Lambda_{\Omega}$ for $i\neq j$. Then $\varphi_{n,k,\gamma,\lambda}(s)$ is a polynomial in $\wp(s)$ of degree $\ell+1$ with the distinct roots $\wp(z_{1}),\ldots,\wp(z_{\ell}),\wp(z)$. Furthermore,
$S_{r}\wp(z_{1},\ldots,z_{\ell},z)=(-1)^{r}\frac{\mu_{n,k,\gamma,\lambda}(\ell+1-r)}{\mu_{n,k,\gamma,\lambda}(\ell+1)}.$
Proof.
Rearranging equation (4), we get
$\sum_{2\nmid n_{i}}\gamma_{i}\wp^{(n_{i})}(z_{j})-\sum_{2\nmid k_{i}}\lambda_{i}\wp^{(k_{i})}(z_{j})=-\sum_{2\mid n_{i}}\gamma_{i}\wp^{(n_{i})}(z_{j})+\sum_{2\mid k_{i}}\lambda_{i}\wp^{(k_{i})}(z_{j}).$
Therefore $\wp(z_{1}),\ldots,\wp(z_{\ell})$ are distinct roots of $\varphi_{n,k,\gamma,\lambda}$. Using the fact that $\wp^{(i)}(-z)=(-1)^{i}\wp^{(i)}(z)$ and rearranging (5), we get
$\sum_{2\nmid n_{i}}\gamma_{i}\wp^{(n_{i})}(z)-\sum_{2\nmid k_{i}}\lambda_{i}\wp^{(k_{i})}(z)=\sum_{2\mid n_{i}}\gamma_{i}\wp^{(n_{i})}(z)-\sum_{2\mid k_{i}}\lambda_{i}\wp^{(k_{i})}(z).$
Thus $\wp(z)$ is a root of $\varphi_{n,k,\gamma,\lambda}$. By Lemma 2.3 and assumption that $\varphi_{n,k,\gamma,\lambda}\not\equiv 0$, this proves that $\varphi_{n,k,\gamma,\lambda}(s)$ is a polynomial in $\wp(s)$ of degree $\ell+1$ with the distinct roots $\wp(z_{1}),\ldots,\wp(z_{\ell}),\wp(z)$. In particular, we see $\mu_{n,k,\gamma,\lambda}(\ell+1)\neq 0$. Therefore we have the factorization
$\varphi_{n,k,\gamma,\lambda}(s)$ $=\mu_{n,k,\gamma,\lambda}(\ell+1)\left(\wp(s)-\wp(z)\right)\prod_{i=1}^{\ell}\wp(s)-\wp(z_{i})$
$=\sum_{r=0}^{\ell+1}\mu_{n,k,\gamma,\lambda}(r)\wp(s)^{r}.$
Comparing the coefficients of $\wp(s)^{r}$, we obtain
$(-1)^{r}\mu_{n,k,\gamma,\lambda}(\ell+1)S_{r}\wp(z_{1},\ldots,z_{\ell},z)=\mu_{n,k,\gamma,\lambda}(\ell+1-r).$
Thus the proof is complete. ∎
Theorem 2.5.
Assume that $S_{r-1}\wp(z_{1},\ldots,z_{\ell})\neq 0$. Then we have
$\wp(z)=\frac{1}{S_{r-1}\wp(z_{1},\ldots,z_{\ell})}\left((-1)^{r}\frac{\mu_{n,k,\gamma,\lambda}(\ell+1-r)}{\mu_{n,k,\gamma,\lambda}(\ell+1)}-S_{r}\wp(z_{1},\ldots,z_{\ell})\right).$
Proof.
The proof follows from trivially by Theorem 2.4 and the identity
$S_{r}\wp(z_{1},\ldots,z_{\ell},z)=\wp(z)S_{r-1}\wp(z_{1},\ldots,z_{\ell})+S_{r}\wp(z_{1},\ldots,z_{\ell}).$
Taking $z_{1},\ldots,z_{\ell}\to s$ in Theorem 2.5, we obtain a multiplication theorem for $\wp(s)$. Comparing Theorem 2.5 for different $r_{1},r_{2}$, we get the following corollary.
Corollary 2.6.
Assume that $S_{r_{1}-1}\wp(z_{1},\ldots,z_{\ell}),S_{r_{2}-1}\wp(z_{1},\ldots,z_{\ell})\neq 0$. Then we have
$(-1)^{r_{1}}S_{r_{2}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{1})-(-1)^{r_{2}}S_{r_{1}-1}\mu_{n,k,\gamma,\lambda}(\ell+1-r_{2})=\left(S_{r_{1}}S_{r_{2}-1}-S_{r_{2}}S_{r_{1}-1}\right)\mu_{n,k,\gamma,\lambda}(\ell+1),$
where $S_{r}=S_{r}\wp(z_{1},\ldots,z_{\ell})$.
We note that $\mu_{n,k,\gamma,\lambda}(r)$ are computable in terms of $\gamma_{i},\lambda_{j}$ by Lemma 2.2 since the expansion of $\wp^{(i)}$ in terms of Weierstrass-$\wp$ is easily obtained by induction and differentiation. Applying Cramer’s rule to the system (4) yields
$\lambda_{j}=\frac{\begin{vmatrix}\wp^{(k_{1})}(z_{1})&\cdots&\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{1})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \wp^{(k_{1})}(z_{2})&\cdots&\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \wp^{(k_{1})}(z_{\ell})&\cdots&\sum_{i=1}^{m}\gamma_{i}\wp^{(n_{i})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\end{vmatrix}}{\begin{vmatrix}\wp^{(k_{1})}(z_{1})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{1})\\ \wp^{(k_{1})}(z_{2})&\wp^{(k_{2})}(z_{2})&\cdots&\wp^{(k_{\ell})}(z_{2})\\ \vdots&\vdots&\vdots&\vdots\\ \wp^{(k_{1})}(z_{\ell})&\wp^{(k_{2})}(z_{\ell})&\cdots&\wp^{(k_{\ell})}(z_{\ell})\end{vmatrix}},$
where only $j$th column of the upper matrix has been replaced.
3 Applications And Discussion
We illustrate our method with the simplest case. We reobtain the classical proof of addition formula Theorem 1.1, as given in [9] and obtain two novel addition and duplication formulas in addition to new identities for elliptic invariants. Let $n_{1}=\gamma_{1}=1$, $(k_{1},k_{2})=(0,-2)$ and $\ell=2$. Then we obtain the system
$\wp^{\prime}(z)=\lambda_{1}\wp(z)+\lambda_{2}$
$\wp^{\prime}(w)=\lambda_{1}\wp(w)+\lambda_{2},$
subject to the condition $\wp(z)\neq\wp(w)$. We then have $\psi(s)=\wp^{\prime}(s)-\lambda_{1}\wp(s)-\lambda_{2}$ and $\varphi(s)=\left(\wp^{\prime}(s)\right)^{2}-(\lambda_{1}\wp(s)+\lambda_{2})^{2}$. By equation (1), we calculate $\mu(r)$ as follows,
$\mu(0)=-\lambda_{2}^{2}-g_{3}=-\left(\frac{\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)}{\wp(z)-\wp(w)}\right)^{2}-g_{3}$
$\mu(1)=-2\lambda_{1}\lambda_{2}-g_{2}=-2\frac{\left(\wp^{\prime}(z)-\wp^{\prime}(w)\right)\left(\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)\right)}{\left(\wp(z)-\wp(w)\right)^{2}}-g_{2}$
$\mu(2)=-\lambda_{1}^{2}=-\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}$
$\mu(3)=4.$
Thus we get the following formulae.
Corollary 3.1.
Assume $z,w,z\pm w\notin\Lambda_{\Omega}$ and $z+w\not\equiv-z,-w\bmod\Lambda_{\Omega}$. Then we have
$\wp(z+w)=\frac{1}{4}\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}-\wp(z)-\wp(w).$ (8)
Furthermore, if $\wp(z)+\wp(w)\neq 0$, then we have
$\wp(z+w)=-\frac{\left(\wp^{\prime}(z)-\wp^{\prime}(w)\right)\left(\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)\right)}{2\left(\wp(z)+\wp(w)\right)\left(\wp(z)-\wp(w)\right)^{2}}-\frac{\wp(z)\wp(w)}{\wp(z)+\wp(w)}-\frac{g_{2}}{4\left(\wp(z)+\wp(w)\right)},$
if $\wp(z)\wp(w)\neq 0$, then we have
$\wp(z+w)=\frac{1}{4\wp(z)\wp(w)}\left(\frac{\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)}{\wp(z)-\wp(w)}\right)^{2}+\frac{g_{3}}{4\wp(z)\wp(w)}.$
Taking the limit $w\to z$, we obtain the classic and two new duplication formulas
$\wp(2z)$ =′′$=-\frac{\wp^{\prime\prime}(z)\left(\wp^{\prime}(z)^{2}-\wp(z)\wp^{\prime\prime}(z)\right)}{4\wp(z)\wp^{\prime}(z)^{2}}-\frac{\wp(z)}{2}-\frac{g_{2}}{8\wp(z)}$
$=\frac{1}{4}\left(\frac{\wp^{\prime}(z)^{2}-\wp(z)\wp^{\prime\prime}(z)}{\wp(z)\wp^{\prime}(z)}\right)^{2}+\frac{g_{3}}{4\wp(z)^{2}},$
which are valid for $\wp(z),\wp^{\prime}(z)\neq 0$, in other words $2z\notin\Lambda$. Above formula can be simplified further and written only in terms of $\wp(z)$ using $\left(\wp^{\prime}\right)^{2}=4\wp^{3}-g_{2}\wp-g_{3}$ and ′′$\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$. Indeed, we have
$\wp(2z)$ $=\frac{16\wp(z)^{4}+8g_{2}\wp(z)^{2}+32g_{3}\wp(z)+g_{2}^{2}}{16\left(4\wp(z)^{3}-g_{2}\wp(z)-g_{3}\right)}$
$=\frac{1}{16}\frac{\left(4\wp(z)^{3}+g_{2}\wp(z)+2g_{3}\right)^{2}}{\wp(z)^{2}\left(4\wp(z)^{3}-g_{2}\wp(z)-g_{3}\right)}+\frac{g_{3}}{4\wp(z)^{2}}.$
Comparing the equations in Corollary 3.1, we have the following identities for the Weierstrassian elliptic invariants $g_{2},g_{3}$.
Corollary 3.2.
Assume $z,w,z\pm w\notin\Lambda_{\Omega}$ and $z+w\not\equiv-z,-w\bmod\Lambda_{\Omega}$. If $\wp(z)+\wp(w)\neq 0$, then we have
$g_{2}$ $=4\left(\wp(z)+\wp(w)\right)^{2}-4\wp(z)\wp(w)-\left(\wp(z)+\wp(w)\right)\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}$
$-2\frac{\left(\wp^{\prime}(z)-\wp^{\prime}(w)\right)\left(\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)\right)}{\left(\wp(z)-\wp(w)\right)^{2}}.$
Furthermore, if $\wp(z)\wp(w)\neq 0$, then we have
$g_{3}=\wp(z)\wp(w)\left(\frac{\wp^{\prime}(z)-\wp^{\prime}(w)}{\wp(z)-\wp(w)}\right)^{2}-4\wp(z)\wp(w)\left(\wp(z)+\wp(w)\right)-\left(\frac{\wp(z)\wp^{\prime}(w)-\wp(w)\wp^{\prime}(z)}{\wp(z)-\wp(w)}\right)^{2}.$
Upon taking $z=\omega_{1}$ and $w=\omega_{2}$ in Corollary 3.2 we get the classic formulae $g_{2}=-4(e_{1}e_{2}+e_{2}e_{3}+e_{3}e_{1})$ and $g_{3}=4e_{1}e_{2}e_{3}$.
We now note several novel three-term addition formulas. Let $n_{1}=2$, $\gamma_{1}=1$, $(k_{1},k_{2},k_{3})=(1,0,-2)$ and $\ell=3$. We have the system
′′$\wp^{\prime\prime}(u)=\lambda_{1}\wp^{\prime}(u)+\lambda_{2}\wp(u)+\lambda_{3}$
′′$\wp^{\prime\prime}(v)=\lambda_{1}\wp^{\prime}(v)+\lambda_{2}\wp(v)+\lambda_{3}$
′′$\wp^{\prime\prime}(w)=\lambda_{1}\wp^{\prime}(w)+\lambda_{2}\wp(w)+\lambda_{3}.$
Using Frobenius-Stickelberger formula (2), condition (3) is equivalent to
$0\neq\begin{vmatrix}1&\wp(u)&\wp^{\prime}(u)\\ 1&\wp(v)&\wp^{\prime}(v)\\ 1&\wp(w)&\wp^{\prime}(w)\end{vmatrix}=2\frac{\sigma(u+v+w)\sigma(u-v)\sigma(v-w)\sigma(w-u)}{\sigma(u)^{3}\sigma(v)^{3}\sigma(w)^{3}}.$
This directly translates to the assumptions that $u,v,w$ are distinct $\bmod\Lambda_{\Omega}$ and $u+v+w\notin\Lambda_{\Omega}$. The determinant in Theorem 2.1 unfortunately vanishes identically by the Frobenius-Stickelberger formula. We get that $\varphi(s)=\lambda_{1}^{2}\wp^{\prime}(s)^{2}-\left(\wp^{\prime\prime}(s)-\lambda_{2}\wp(s)-\lambda_{3}\right)^{2}$. Utilizing (1) and its corollary ′′$\wp^{\prime\prime}=6\wp^{2}-g_{2}/2$, we calculate $\mu(r)$ as follows,
$\mu(0)=-\lambda_{1}^{2}g_{3}-\lambda_{3}^{2}-\lambda_{3}g_{2}-\frac{g_{2}^{2}}{4}$
$\mu(1)=-\lambda_{1}^{2}g_{2}-2\lambda_{2}\lambda_{3}-\lambda_{2}g_{2}$
$\mu(2)=-\lambda_{2}^{2}+12\lambda_{3}+6g_{2}$
$\mu(3)=4\lambda_{1}^{2}+12\lambda_{2}$
$\mu(4)=-36.$
By Cramer’s rule, $\lambda_{1},\lambda_{2}$and $\lambda_{3}$ are given by
$\lambda_{1}=\frac{\sum_{cyc}\wp^{\prime\prime}(u)\wp(v)-\wp(u)\wp^{\prime\prime}(v)}{\sum_{cyc}\wp^{\prime}(u)\wp(v)-\wp(u)\wp^{\prime}(v)}$
$\lambda_{2}=\frac{\sum_{cyc}\wp^{\prime}(u)\wp^{\prime\prime}(v)-\wp^{\prime\prime}(u)\wp^{\prime}(v)}{\sum_{cyc}\wp^{\prime}(u)\wp(v)-\wp(u)\wp^{\prime}(v)}$
$\lambda_{3}=\frac{\sum_{cyc}\wp(u)(\wp^{\prime\prime}(v)\wp^{\prime}(w)-\wp^{\prime}(v)\wp^{\prime\prime}(w))}{\sum_{cyc}\wp^{\prime}(u)\wp(v)-\wp(u)\wp^{\prime}(v)},$
where summations are cyclic in terms of $u,v,w$. Thus we have the following addition theorems.
Corollary 3.3.
Let $u,v,w$ be distict $\bmod\Lambda_{\Omega}$, $u,v,w,u+v+w\notin\Lambda_{\Omega}$ and $u+v+w\not\equiv-u,-v,-w\bmod\Lambda_{\Omega}$. Then we have
$\wp(u+v+w)=\frac{\lambda_{1}^{2}}{9}+\frac{\lambda_{2}}{3}-\wp(u)-\wp(v)-\wp(w).$ (9)
If $\wp(u)+\wp(v)+\wp(w)\neq 0$, then
$\wp(u+v+w)=\frac{\lambda_{2}^{2}-12\lambda_{3}-6g_{2}}{36(\wp(u)+\wp(v)+\wp(w))}-\frac{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}{\wp(u)+\wp(v)+\wp(w)}$
If $\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)\neq 0$, then
$\wp(u+v+w)=-\frac{\lambda_{1}^{2}g_{2}+2\lambda_{2}\lambda_{3}+\lambda_{2}g_{2}}{36(\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u))}-\frac{\wp(u)\wp(v)\wp(w)}{\wp(u)\wp(v)+\wp(v)\wp(w)+\wp(w)\wp(u)}.$
If $\wp(u)\wp(v)\wp(w)\neq 0$, then
$\wp(u+v+w)=\frac{4\lambda_{1}^{2}g_{3}+4\lambda_{3}^{2}+4\lambda_{3}g_{2}+g_{2}^{2}}{144\wp(u)\wp(v)\wp(w)}.$
Now we take the limits $v,w\to u=z$ to obtain triplication formula for $\wp(z)$. After a lengthy computation of limits we obtain
$\lim_{u,v,w\to z}\lambda_{1}=\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}$
$\lim_{u,v,w\to z}\lambda_{2}=\frac{\wp^{\prime\prime}(z)\wp^{(4)}(z)-\wp^{\prime\prime\prime}(z)^{2}}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}$
$\lim_{u,v,w\to z}\lambda_{3}=\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}$
Hence subject to the condition that $z,3z,4z\notin\Lambda_{\Omega}$, we have the following triplication formula,
$\wp(3z)=\frac{1}{9}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}+\frac{1}{3}\frac{\wp^{\prime\prime}(z)\wp^{(4)}(z)-\wp^{\prime\prime\prime}(z)^{2}}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}-3\wp(z).$ (10)
If we assume that in addition $\wp(z)\neq 0$, we get the more complicated ones
$\wp(3z)$ $=\frac{1}{108\wp(z)}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}-\frac{g_{2}}{18\wp(z)}-\wp(z)$
$-\frac{1}{9\wp(z)}\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}$
and
$\wp(3z)$ $=-\frac{\wp(z)}{3}-\frac{g_{2}}{108\wp(z)^{2}}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}-\frac{1}{108\wp(z)^{2}}\left(\frac{\wp^{\prime\prime}(z)\wp^{(4)}(z)-\wp^{\prime\prime\prime}(z)^{2}}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)$
×$\times\left(g_{2}+2\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)$
and
$\wp(3z)$ $=\frac{g_{2}^{2}}{144\wp(z)^{3}}+\frac{g_{3}}{36\wp(z)^{3}}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}$
$+\frac{g_{2}}{36\wp(z)^{3}}\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}$
$+\frac{1}{36\wp(z)^{3}}\left(\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}.$
We obviously recommend the formula (10) for practical use. Above equations can be written only in terms of $\wp(z)$ utilizing the following
′′′$\wp^{\prime\prime\prime}=12\wp^{\prime}\wp\qquad\text{and}\qquad\wp^{(4)}=120\wp^{3}-18g_{2}\wp-12g_{3}.$
Substituting these in the definitions of $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ we obtain
$\lim_{u,v,w\to z}\lambda_{1}=\frac{48\wp^{\prime}\left(4\wp^{3}-g_{2}\wp-g_{3}\right)}{48\wp^{4}-24g_{2}\wp^{2}-48g_{3}\wp-g_{2}^{2}}$
$\lim_{u,v,w\to z}\lambda_{2}=-\frac{576\wp^{5}-96g_{2}\wp^{3}+288g_{3}\wp^{2}+36g_{2}^{2}\wp+24g_{2}g_{3}}{48\wp^{4}-24g_{2}\wp^{2}-48g_{3}\wp-g_{2}^{2}}$
$\lim_{u,v,w\to z}\lambda_{3}=\frac{192\wp^{6}+240g_{2}\wp^{4}+768g_{3}\wp^{3}-12g_{2}^{2}\wp^{2}-96g_{2}g_{3}\wp-96g_{3}^{2}+g_{2}^{3}}{96\wp^{4}-48g_{2}\wp^{2}-96g_{3}\wp-2g_{2}^{2}},$
where $\wp=\wp(z)$. For example equation (10) is equivalent to
$\left(48\wp^{4}-24g_{2}\wp^{2}-48g_{3}\wp-g_{2}^{2}\right)^{2}\wp(3z)$ $=256\wp^{9}+768g_{2}\wp^{7}+6144g_{3}\wp^{6}+480g_{2}^{2}\wp^{5}-384g_{2}g_{3}\wp^{4}$
$+\left(768g_{3}^{2}-144g_{2}^{3}\right)\wp^{3}-192g_{2}^{2}g_{3}\wp^{2}+\left(9g_{2}^{4}-384g_{2}g_{3}^{2}\right)\wp$
$+8g_{2}^{3}g_{3}-256g_{3}^{3}$
We remark that one might obtain simpler formulas such as (8) and (9) for $\wp(z_{1}+\cdots+z_{\ell})$ by considering the system
$\wp^{(\ell-1)}(z_{j})=\lambda_{1}\wp^{(\ell-2)}(z_{j})+\lambda_{2}\wp^{(\ell-3)}(z_{j})+\cdots+\lambda_{\ell-1}\wp(z_{j})+\lambda_{\ell}.$
This corresponds to the case $n_{1}=\ell-1$, $\gamma_{1}=1$, $(k_{1},\cdots,k_{\ell-1},k_{\ell})=(\ell-2,\cdots,0,-2)$. Similarly, the determinant in condition (3) can be calculated by Frobenius-Stickelberger formula and is equivalent to the statement that $z_{i}$ are distinct $\bmod\Lambda_{\Omega}$ with $z\notin\Lambda$. Taking the limits $z_{1},\cdots,z_{\ell}\to s$, we get the $\ell$-multiplication formula for $\wp(s)$.